Question 22

For a two-factor ANOVA, the degrees of freedom values are as follows:

dfbetween = 5 dfA = 1 dfB = 2

What is the value of dfAxB?

A 2
B 1
C 5
D 3

SHOW WORK
Answer #1

For two-factor ANOVA, the degrees of freedom values are calculated as

df A * B = df A * df B

= 2 * 1

df A * B = 2